### v2

# Diff-MN: Diffusion Parameterized MoE-NCDE for Continuous Time Series Generation with Irregular Observations

Abstract
Time series generation (TSG) is widely used across domains, yet most existing methods assume regular sampling and fixed output resolutions. These assumptions are often violated in practice, where observations are irregular and sparse, while downstream applications require continuous and high-resolution TS. Although Neural Controlled Differential Equation (NCDE) is promising for modeling irregular TS, it is constrained by a single dynamics function, tightly coupled optimization, and limited ability to adapt learned dynamics to newly generated samples from the generative model. We propose Diff-MN, a continuous TSG framework that enhances NCDE with a Mixture-of-Experts (MoE) dynamics function and a decoupled architectural design for dynamics-focused training. To further enable NCDE to generalize to newly generated samples, Diff-MN employs a diffusion model to parameterize the NCDE temporal dynamics parameters (MoE weights), i.e., jointly learn the distribution of TS data and MoE weights. This design allows sample-specific NCDE parameters to be generated for continuous TS generation. Experiments on ten public and synthetic datasets demonstrate that Diff-MN consistently outperforms strong baselines on both irregular-to-regular and irregular-to-continuous TSG tasks. The code is available at the link https://github.com/microsoft/TimeCraft/tree/main/Diff-MN.
Keywords Time Series Generation Diffusion Models Mixture of Experts Neural Controlled Differential Equations Irregular Observations
1 Introduction
Time series generation (TSG) is essential across many real-world domains, including healthcare (Deng et al., 2025a, b), finance (Zhang et al., 2025a; Vasiliu et al., 2024; Zhang et al., 2024b), manufacturing and energy systems (Zhang et al., 2025b; Fuest et al., 2025; Zhang et al., 2025c). It supports data augmentation (Tang et al., 2025), privacy-preserving data sharing (Tian et al., 2024), and high-fidelity simulation for decision-making (Turowski et al., 2024). Despite substantial progress, most TSG methods assume regularly sampled TS, treating both training and generated data as discrete observations at fixed intervals. As shown in Fig. 1(a), existing approaches typically train on uniformly observed sequences and generate outputs at fixed resolutions (Mushunje et al., 2023; Li, 2023; Huang et al., 2025; Ge et al., 2025). While a few methods allow irregular training data, they still generate sequences at fixed intervals, as illustrated in Fig. 1(b) (Nikitin et al., 2024; Ramponi et al., 2018). This setting mismatches real-world scenarios, where observations are often irregular and sparse, yet downstream tasks require continuous and high-resolution TS. For example, in ICU monitoring, electronic health records are irregular and sparse, while applications such as patient trajectory simulation or risk forecasting require dense sequences (Tian et al., 2024; Li et al., 2025a). Consequently, continuous TSG in Fig. 1(c) from irregular observations remains an under-explored but critical challenge.
Continuous-time modeling and irregular observations have been widely studied, leading to several approaches. Among them, Neural Ordinary Differential Equations (NODEs) and their extension, Neural Controlled Differential Equations (NCDE), are particularly popular. By combining continuous-time differential equations with learnable neural networks, NCDE-based methods effectively model temporal dynamics and handle irregular observations. Recent work has applied them to TSG (Lim et al., 2023; Naiman et al., 2024), typically following a two-stage paradigm: irregularly sampled TS are first transformed into regularly sampled ones using NCDE or their variants, and standard TSG models such as GANs (Jeon et al., 2022) and VAEs (Naiman et al., 2024) are then applied to generate data on regular time grids.
However, on the one hand, existing methods mainly use NCDE to process irregular inputs and then generate regularly sampled TS (Fig. 1(b)), with limited exploration of continuous TSG and evaluation (Fig. 1(c)). On the other hand, they directly adopt the standard NCDE without further improvement. Although standard NCDE are effective for modeling irregular TS, they rely on a single dynamics function, lack dynamics-focused optimization design, and cannot generalize learned dynamics to newly generated samples from a generative model. As a result, there remains substantial room to improve standard NCDE for irregular TS modeling and high-quality continuous-time generation.
In this study, we propose Diff-MN, a framework for continuous TSG with irregular observations. Our work differs from existing studies in the following aspects:
1.
We are the first to formally study continuous TSG and evaluate it under the generated high-frequency TS data from two perspectives: TS forecasting performance and analytical solutions of cubic polynomials.
2.
To enable effective continuous TSG, we improve standard NCDE in two key ways. (i) We replace the single dynamics function with a mixture-of-experts (MoE) dynamics function, where different experts capture distinct temporal patterns and are aggregated to model complex dynamics. The learned MoE weights explicitly encode temporal dynamics, providing a basis for using the diffusion model to parameterize them, thereby generalizing to newly generated samples. (ii) We propose a dynamics-focused, decoupled optimization strategy for NCDE. Specifically, we replace the State Initialization Network (SIN) and Readout Network (RN) with a pre-trained channel-wise autoencoder, which warm-starts training and allows the model to focus on optimizing the dynamics function.
3.
Based on (i) and (ii), the MoE weights learned by the trained NCDE encode temporal dynamics that are specific to the training samples. When new samples are generated, their temporal dynamics may differ, leading to a mismatch with these sample-specific weights. To address this, we propose to parameterize the MoE weights for each new sample. Specifically, we jointly model TS data and their corresponding MoE weights with a diffusion model to learn their joint distribution. As a result, for each newly generated sample, matched MoE weights that reflect the true dynamics are generated and substituted back into the pre-trained NCDE, enabling more accurate continuous time series generation.
Our contributions are as follows:
We are the first to formally study continuous TSG and evaluate it on generated high-frequency time series, filling an important research gap.
For better continuous TSG, we first propose a decoupled Mixture-of-Experts Neural CDE (MoE-NCDE), which enhances standard NCDE with the MoE dynamic functions and dynamics-focused optimization design. This enables NCDE to better capture complex and dynamically changing temporal patterns from irregular observations.
We further propose Diff-MN, a continuous TSG framework that jointly models TS data and MoE-NCDE temporal dynamics parameters (MoE weights) using a diffusion model. This design parameterizes the MoE weights and allows sample-specific weights to be generated and substituted back into a pre-trained MoE-NCDE, enabling accurate continuous TSG for newly generated samples.
We conduct extensive experiments on four popular time-series forecasting datasets, four medical ECG datasets, and two synthetic datasets. Both quantitative and qualitative results demonstrate that Diff-MN consistently outperforms strong baselines on both irregular-to-regular and irregular-to-continuous generation tasks.
2 Related work
The most relevant methods are NCDE-based generators, e.g., KOVAE (Jeon et al., 2022) and GTGAN (Naiman et al., 2024), which build on standard NCDE and conduct irregular-to-regular generations. Some approaches also attempt to use attention-style mechanisms to capture the relationship between the observed sparse values and the dense time grids, learning the time dependencies and thereby enabling prediction at arbitrary time-steps (Shukla and Marlin, ; Yalavarthi et al., 2025). Unlike existing methods, we are the first to formally study the continuous TSG task and propose the Diff-MN framework, which introduces a mixture-of-experts dynamics function, a decoupled NCDE architectural design, and diffusion-based parameterization of MoE weights to generate sample-specific NCDE dynamics for each newly generated TS. Complete related work about TSG models for regular and irregular data is provided in the Appendix A.
3 Method
Overview. Our method is shown in Figure 2. We next detail the training and generation process of Diff-MN.
Training: First, we train the decoupled MoE NCDE on irregular observations, learning shared State Initialization Network (SIN), Readout Network (RN), and dynamic-function neural networks, along with sample-specific MoE weights that encode each instance’s temporal dynamics. These components are then used to impute missing values. Next, to parameterize the MoE weights for generalizing to new samples, we jointly train a diffusion model on the imputed data and their corresponding MoE weights. Consequently, the diffusion model generates new samples together with matched MoE weights that reflect their temporal dynamics, enabling better continuous TSG.
Generation: The newly generated samples and MoE weights from the diffusion model are fed back into pretrained MoE-NCDE for continuous TSG, producing high-frequency, longer, and more informative TS.
3.1 Preliminary
Notation. Suppose we have observed multivariate TS. Each TS , for , is a continuous multivariate function defined over its time interval, where is the number of channels/variables. In practice, we observe each multivariate TS at discrete and irregular time points , where . The observed dataset is:
Problem Definition. Given dataset , the goal is to learn a generative model over discrete and irregular TS for irregular or continuous generation, i.e., our goal is to learn a model capable of sampling new sequences . The generated dataset is denoted as where denotes arbitrary time step of -th sample.
Neural Controlled Differential Equation. Neural CDE (Kidger et al., 2020) extends Neural ODE (Chen et al., 2018) by modeling latent dynamics driven by a continuous control path. In continuous TSG, values generated between observations must align with surrounding temporal dynamics to avoid deviating from the true system distribution. Theoretically, we use NCDE instead of NODE because NCDE conditions the dynamics on irregular observations via a control path, enabling more accurate interpolation, whereas NODE may drift without observation guidance. The details of why NCDE is better in our task from both theoretical and experimental perspectives can be found in Appendix B.1.
The NCDE is defined as:
where is a learnable dynamic-function neural network, and is a continuous interpolation (e.g., a natural cubic spline) of the discrete observations . is the initial latent state mapped from based on the network SIN.
Notably, although the control path is smooth due to spline interpolation, the latent trajectory can still be non-smooth because its evolution is governed by the nonlinear neural network and the mixture-of-experts, which together allow expressive, non-smooth, and highly diverse temporal patterns. The proof from both theoretical and experimental perspectives can be found in Appendix B.2.
Finally, CDESolver to solve NCDE is formalized as:
Its output is the latent state at arbitrary time points . Then, the network RN subsequently maps this hidden representation back to the original channel space to produce the predicted TS values.
3.2 Modeling diverse dynamics with dense MoE
Real-world TS exhibit diverse temporal patterns that a single dynamic function often struggles to capture. Using a Mixture of Experts (MoE), the complex temporal dynamics are first decomposed, allowing each expert to focus on the patterns it is best suited to learn, such as local or global trends. This makes the overall dynamics easier to learn while enhancing both expressiveness and generalization. By combining their outputs, the MoE-NCDE produces richer and more flexible temporal representations. This is strongly supported by Figure 6 in the experimental section.
Notably, we use a dense MoE rather than a sparse one, prioritizing stability over efficiency. Dense MoE fully engages all experts at each step, enabling stable training and better learning of complex temporal dynamics in our task, which involves limited training data and low-parameter models.
The MoE dynamics consist of a MLP (Multilayer Perceptron) based router and dynamic functions (each parameterized as a neural network), where . The router takes (potentially irregular) observations as input and outputs a routing score for each expert. Specifically, we first impute the irregular TS using cubic interpolation and then feed them into to obtain all experts’ scores for the -th sample. The router is implemented as a simple MLP. A softmax function is then applied over the scores to produce a dense routing distribution. Given , the final output is computed as a weighted combination of all expert functions:
Finally, substituting into the standard NCDE formulas (Eq. 2 and Eq. 3) yields the MoE-NCDE.
Notably, the learned MoE weights explicitly encode the temporal dynamics of each training sample. This property provides a foundation for parameterizing the MoE via a diffusion model. Consequently, for newly generated samples, the diffusion model can produce adapted MoE weights that capture their temporal dynamics, enabling customized and effective continuous TSG. A detailed description will be provided in Section 3.4.
3.3 Decoupled optimization design for MoE-dynamics-focused training
As shown in Figure 2(a), training a standard NCDE requires jointly optimizing the State Initialization Network (SIN), the Readout Network (RN), and the dynamic function network. Joint training can cause conflicts among modules and make the dynamic function network, which is the most critical component, difficult to learn effectively, especially when SIN and RN are poorly initialized.
To address this, we pretrain a channel-wise autoencoder on the irregular data and replace SIN and RN with the pretrained Channel-wise Encoder (CE) and Decoder (CD), keeping them frozen. This decoupled optimization (termed Decoupled Design) enables training to focus on the optimization of MoE dynamic functions while using CE and CD for a warm start, providing effective gradients and improving training stability. This is feasible because the pretrained CE and CD have the same input and output dimensions as SIN and RN, allowing direct substitution. The training of the CE and CD is shown in Appendix Algorithm 1 and overall MoE-dynamics-focused training in Appendix Algorithm 2.
3.4 Diffusion parameterized MoE-NCDE for continuous TSG
Building on the MoE-NCDE trained on a batch of samples , it can generate continuous time sequences by predicting values at any consistent with contextual time information. However, the learned MoE weights, which encode temporal dynamics, are specific to the training samples. New samples generated via a diffusion model have different temporal dynamics, rendering the original MoE weights incompatible. To address this, we propose parameterizing the MoE weights with a diffusion model, enabling newly generated samples to obtain adapted MoE weights that capture their temporal dynamics and allow customized, effective continuous TSG.
Specifically, the standard diffusion model requires regular data, but our training samples are irregular. We first use the pretrained MoE-NCDE to impute each sample into a regular sequence. The regularized samples, along with their corresponding MoE weights , are then fed into the diffusion model to learn the joint distribution of temporal patterns and MoE weights. This parameterizes the MoE weights, allowing new samples to receive adapted weights that describe their temporal dynamics. The training process is shown in the Appendix algorithm 3.
Finally, we input into the pre-trained MoE-NCDE and perform continuous TSG for each sample using its generated MoE weights and CDESolver. The process is shown in Appendix Algorithm 4.
4 Experiments
The experiment aims to answer the following questions:
4.1 Experimental settings
Datasets. The datasets include: (1) four public TSG datasets, Sines, Stocks, Energy, and MuJoCo; (2) ECG-related medical TS datasets from the UCR archive with diverse temporal patterns, including ECG200, ECG5K, ECGFD, and TLECG, with 2, 5,2 and 2 classes respectively; and (3) synthetic datasets, consisting of integrated signal waves and polynomial coefficient fitting data. For irregular modeling, we randomly discard 30%, 50%, and 70% of temporal observations to construct irregularly sampled datasets (Naiman et al., 2024). Details are provided in Appendix D.1.
Baselines. We compare our method with several representative advanced TSG methods, including irregular TSG methods: KO-VAE (Naiman et al., 2024), GTGAN (Jeon et al., 2022), ProFITi Yalavarthi et al. (2025) and HeTVAE (Shukla and Marlin, ), as well as regular TSG methods: TimeGAN (Yoon et al., 2019), TimeVAE (Desai et al., 2021), and diffusion model (Huang et al., 2025). For regular TSG methods, we utilize standard NCDE to impute missing values before generation.
Implementation details. We conduct experiments using publicly available baseline code and train all models to convergence with recommended settings. In our method, the number of experts is fixed at four. More parameter settings, e.g., the diffusion model, are provided in the Appendix D.2.
For the four public datasets and two synthetic datasets, we construct training and evaluation samples with three sequence lengths (12, 24, and 36) and report averaged metrics. For the medical ECG classification datasets, samples are officially preprocessed, and default parameters are used.
Evaluation metrics. Following previous works (Naiman et al., 2024; Huang et al., 2025), we apply the following metrics for evaluating generation performance: (1) Discriminative score (DS), (2) Marginal distribution difference (MDD), (3) Kullback-Leibler divergence (KL), (4) Membership Inference Risk (MIR) metric (Tian et al., 2024) for privacy protection evaluation.
4.2 Performance of irregular-to-regular TSG
Quantitative evaluation. We first evaluate our model on the irregular TSG task. As shown in Tables 1 and 2, our method consistently outperforms baselines on both popular TSG and medical datasets across various missing rate settings. Although baselines handle irregular data via NCDE, their performance is limited by weak dynamic modeling and coupled optimization design. In contrast, our MoE-NCDE leverages a mixture-of-experts function and a dynamics-focused optimization strategy with a pre-trained channel-wise autoencoder, enabling more effective learning of temporal dynamics for irregular TS.
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| DS | Ours | 0.105 | 0.142 | 0.422 | 0.293 | 0.128 | 0.137 | 0.487 | 0.375 | 0.182 | 0.106 | 0.497 | 0.393 |
| KoVAE | 0.142 | 0.225 | 0.476 | 0.372 | 0.171 | 0.187 | 0.500 | 0.397 | 0.228 | 0.213 | 0.500 | 0.425 | |
| GT-GAN | 0.302 | 0.321 | 0.499 | 0.489 | 0.416 | 0.326 | 0.500 | 0.494 | 0.341 | 0.292 | 0.499 | 0.492 | |
| TimeGAN-NCDE | 0.455 | 0.365 | 0.499 | 0.49 | 0.427 | 0.442 | 0.499 | 0.499 | 0.44 | 0.477 | 0.500 | 0.499 | |
| TimeVAE-NCDE | 0.234 | 0.470 | 0.498 | 0.375 | 0.308 | 0.485 | 0.496 | 0.463 | 0.426 | 0.492 | 0.498 | 0.476 | |
| Diffusion-NCDE | 0.200 | 0.441 | 0.460 | 0.355 | 0.330 | 0.475 | 0.489 | 0.427 | 0.421 | 0.491 | 0.497 | 0.473 | |
| ProFITi | 0.307 | 0.454 | 0.495 | 0.492 | 0.344 | 0.483 | 0.499 | 0.497 | 0.392 | 0.489 | 0.5 | 0.497 | |
| HeTVAE | 0.491 | 0.328 | 0.497 | 0.499 | 0.5 | 0.443 | 0.5 | 0.499 | 0.5 | 0.487 | 0.5 | 0.5 | |
| MDD | Ours | 0.953 | 0.25 | 0.270 | 0.347 | 1.093 | 0.281 | 0.252 | 0.318 | 1.308 | 0.299 | 0.279 | 0.297 |
| KoVAE | 5.134 | 0.709 | 0.347 | 0.374 | 5.584 | 0.662 | 0.368 | 0.419 | 6.393 | 0.624 | 0.377 | 0.619 | |
| GT-GAN | 2.644 | 0.630 | 0.567 | 0.558 | 2.978 | 0.589 | 0.473 | 0.596 | 2.559 | 0.603 | 0.532 | 0.627 | |
| TimeGAN-NCDE | 3.477 | 1.111 | 1.028 | 1.182 | 3.739 | 1.605 | 1.343 | 1.535 | 3.981 | 1.629 | 1.559 | 1.704 | |
| TimeVAE-NCDE | 4.767 | 0.597 | 0.444 | 0.338 | 3.761 | 0.726 | 0.533 | 0.443 | 3.397 | 0.843 | 0.604 | 0.565 | |
| Diffusion-NCDE | 1.462 | 0.341 | 0.292 | 0.371 | 1.498 | 0.476 | 0.351 | 0.37 | 1.997 | 0.743 | 0.513 | 0.487 | |
| ProFITi | 1.163 | 0.297 | 0.255 | 0.295 | 1.236 | 0.408 | 0.35 | 0.353 | 1.443 | 0.445 | 0.442 | 0.448 | |
| HeTVAE | 3.158 | 0.599 | 0.625 | 0.654 | 3.449 | 0.773 | 1.095 | 1.093 | 3.967 | 1.032 | 1.323 | 1.398 | |
| KL | Ours | 0.013 | 0.074 | 0.020 | 0.021 | 0.023 | 0.094 | 0.022 | 0.009 | 0.033 | 0.091 | 0.017 | 0.014 |
| KoVAE | 4.172 | 1.014 | 0.065 | 0.130 | 4.719 | 0.859 | 0.067 | 0.198 | 5.521 | 0.868 | 0.08 | 0.425 | |
| GT-GAN | 0.099 | 0.203 | 0.056 | 0.021 | 0.153 | 0.181 | 0.051 | 0.028 | 0.163 | 0.223 | 0.057 | 0.048 | |
| TimeGAN-NCDE | 4.963 | 1.742 | 0.259 | 0.560 | 2.954 | 6.776 | 0.585 | 1.153 | 12.478 | 11.63 | 0.995 | 1.810 | |
| TimeVAE-NCDE | 1.999 | 0.224 | 0.088 | 0.075 | 1.487 | 0.259 | 0.101 | 0.118 | 1.196 | 0.345 | 0.136 | 0.172 | |
| Diffusion-NCDE | 0.056 | 0.085 | 0.065 | 0.045 | 0.070 | 0.094 | 0.058 | 0.058 | 0.095 | 0.105 | 0.087 | 0.078 | |
| ProFITi | 0.079 | 0.074 | 0.022 | 0.015 | 0.112 | 0.121 | 0.033 | 0.015 | 0.138 | 0.132 | 0.045 | 0.031 | |
| HeTVAE | 1.178 | 1.085 | 0.248 | 0.258 | 4.236 | 1.4 | 0.917 | 0.996 | 2.488 | 1.612 | 1.578 | 2.202 | |
| 30% | 50% | 70% | |||||||||||
| ECG200 | ECG5k | ECGFD | TLECG | ECG200 | ECG5k | ECGFD | TLECG | ECG200 | ECG5k | ECGFD | TLECG | ||
| DS | Ours | 0.188 | 0.182 | 0.120 | 0.240 | 0.250 | 0.326 | 0.070 | 0.160 | 0.390 | 0.424 | 0.140 | 0.340 |
| KoVAE | 0.500 | 0.500 | 0.420 | 0.400 | 0.485 | 0.498 | 0.410 | 0.440 | 0.485 | 0.500 | 0.380 | 0.400 | |
| GT-GAN | 0.467 | 0.493 | 0.450 | 0.490 | 0.457 | 0.495 | 0.380 | 0.470 | 0.475 | 0.497 | 0.430 | 0.490 | |
| TimeGAN-NCDE | 0.495 | 0.500 | 0.410 | 0.500 | 0.482 | 0.478 | 0.420 | 0.500 | 0.490 | 0.500 | 0.420 | 0.490 | |
| TimeVAE-NCDE | 0.440 | 0.371 | 0.350 | 0.430 | 0.427 | 0.370 | 0.380 | 0.440 | 0.415 | 0.400 | 0.400 | 0.450 | |
| Diffusion-NCDE | 0.458 | 0.334 | 0.390 | 0.430 | 0.460 | 0.396 | 0.430 | 0.430 | 0.458 | 0.429 | 0.380 | 0.400 | |
| ProFITi | 0.48 | 0.496 | 0.3 | 0.45 | 0.462 | 0.497 | 0.38 | 0.45 | 0.48 | 0.498 | 0.36 | 0.46 | |
| HeTVAE | 0.5 | 0.496 | 0.42 | 0.5 | 0.5 | 0.498 | 0.43 | 0.5 | 0.5 | 0.5 | 0.43 | 0.5 | |
| MDD | Ours | 0.25 | 0.09 | 0.85 | 1.016 | 0.247 | 0.109 | 0.794 | 0.972 | 0.253 | 0.108 | 0.833 | 1.064 |
| KoVAE | 0.551 | 0.364 | 1.588 | 1.467 | 0.565 | 0.359 | 1.581 | 1.439 | 0.523 | 0.357 | 1.604 | 1.475 | |
| GT-GAN | 0.495 | 0.409 | 0.97 | 1.158 | 0.506 | 0.437 | 0.968 | 1.091 | 0.556 | 0.296 | 0.99 | 1.177 | |
| TimeGAN-NCDE | 0.807 | 0.592 | 1.609 | 1.656 | 0.804 | 0.573 | 1.683 | 1.679 | 0.806 | 0.558 | 1.608 | 1.681 | |
| TimeVAE-NCDE | 0.417 | 0.173 | 1.026 | 1.136 | 0.37 | 0.175 | 0.970 | 1.265 | 0.363 | 0.154 | 0.939 | 1.131 | |
| Diffusion-NCDE | 0.389 | 0.187 | 0.945 | 1.125 | 0.369 | 0.190 | 0.973 | 1.082 | 0.328 | 0.198 | 0.958 | 1.130 | |
| ProFITi | 0.31 | 0.185 | 0.989 | 1.074 | 0.298 | 0.196 | 1.006 | 1.115 | 0.325 | 0.247 | 1.003 | 1.138 | |
| HeTVAE | 0.807 | 0.492 | 1.51 | 1.608 | 0.806 | 0.461 | 1.538 | 1.576 | 0.81 | 0.569 | 1.532 | 1.594 | |
| KL | Ours | 0.053 | 0.014 | 0.052 | 0.129 | 0.09 | 0.021 | 0.099 | 0.073 | 0.113 | 0.018 | 0.04 | 0.153 |
| KoVAE | 8.645 | 7.584 | 10.158 | 9.482 | 8.693 | 7.537 | 10.112 | 9.440 | 8.692 | 7.537 | 10.144 | 9.477 | |
| GT-GAN | 5.129 | 3.305 | 2.004 | 5.314 | 5.013 | 3.592 | 2.543 | 7.912 | 7.126 | 2.672 | 3.754 | 12.539 | |
| TimeGAN-NCDE | 17.57 | 16.615 | 13.165 | 17.182 | 17.56 | 14.408 | 13.154 | 17.191 | 17.57 | 9.431 | 13.165 | 17.191 | |
| TimeVAE-NCDE | 0.917 | 0.113 | 2.236 | 2.397 | 0.641 | 0.096 | 2.106 | 3.005 | 0.731 | 0.073 | 2.103 | 1.386 | |
| Diffusion-NCDE | 0.639 | 0.095 | 1.867 | 0.788 | 0.87 | 0.125 | 1.156 | 0.632 | 0.365 | 0.136 | 1.117 | 1.336 | |
| ProFITi | 0.085 | 0.053 | 0.204 | 0.245 | 0.073 | 0.044 | 0.262 | 0.346 | 0.149 | 0.066 | 0.395 | 1.188 | |
| HeTVAE | 17.216 | 8.574 | 8.218 | 16.49 | 17.199 | 6.251 | 13.445 | 16.41 | 17.193 | 10.316 | 13.445 | 16.458 | |
Qualitative evaluation. We also qualitatively evaluate the similarity between real and generated sequences using two visualizations: (i) t-SNE projections and (ii) kernel density estimation of probability density functions (PDFs). As shown in Figure 3, the t-SNE plots (top) exhibit strong overlap between real and synthetic samples, while the PDFs (bottom) show closer alignment with real data compared to KO-VAE and GT-GAN.
Privacy protection analysis. This experiment evaluates whether synthetic data generated by different models leaks information from real data, with results reported in Appendix Table 9. Our method not only achieves superior fidelity in DS, KL, and MDD but also exhibits the lowest privacy leakage risk according to MIR. This is due to the MoE design, which promotes diverse generated samples rather than mere replication of the original data.
4.3 Performance of irregular-to-continuous TSG
Refined continuous TSG on the downstream task. In this experiment, we evaluate continuous TSG with irregular observations using a multivariate prediction task, which is conducted on the newly generated samples of each method. Baselines predict the last time step from length data. For Diff-MN, we apply MoE-NCDE for newly generated samples and MoE weights from the diffusion model. For KOVAE and GTGAN, we apply the trained standard NCDE to their newly generated samples for Continuous TSG. Continuous TSG generates new points between each time step, creating a refined sample that roughly doubles the length of the historical window. Following previous downstream evaluations Naiman et al. (2024); Jeon et al. (2022), the refined sample is then used to train the Gated Recurrent Unit (GRU) to predict the same last step. We set and report average MSE and MAE.
Table 3 shows that KOVAE-NCDE and GTGAN-NCDE perform worse than their non-continuous counterparts, indicating that standard NCDE for continuous generation is ineffective. In contrast, Ours- outperforms its non-continuous version, improving forecasting performance. This is due to our diffusion parameterized MoE-NCDE, which generates MoE weights capturing the temporal dynamics of the newly generated samples, enabling better continuous generation. Visualizations of continuous TS are in Appendix Figures 10 to 12.
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| MSE | Ours- | 0.026 | 0.002 | 0.013 | 0.029 | 0.026 | 0.002 | 0.013 | 0.029 | 0.028 | 0.002 | 0.013 | 0.029 |
| Ours | 0.049 | 0.005 | 0.014 | 0.031 | 0.05 | 0.005 | 0.014 | 0.030 | 0.054 | 0.005 | 0.014 | 0.030 | |
| KOVAE-NCDE | 0.076 | 0.041 | 0.023 | 0.030 | 0.076 | 0.036 | 0.021 | 0.032 | 0.079 | 0.0300 | 0.020 | 0.033 | |
| KOVAE | 0.044 | 0.025 | 0.014 | 0.030 | 0.044 | 0.019 | 0.014 | 0.031 | 0.045 | 0.014 | 0.014 | 0.032 | |
| GTGAN-NCDE | 0.089 | 0.021 | 0.027 | 0.053 | 0.076 | 0.018 | 0.026 | 0.053 | 0.070 | 0.021 | 0.028 | 0.058 | |
| GTGAN | 0.044 | 0.014 | 0.019 | 0.042 | 0.051 | 0.008 | 0.017 | 0.047 | 0.046 | 0.011 | 0.019 | 0.047 | |
| MAE | Ours- | 0.125 | 0.025 | 0.075 | 0.129 | 0.126 | 0.026 | 0.073 | 0.131 | 0.128 | 0.025 | 0.075 | 0.131 |
| Ours | 0.174 | 0.036 | 0.075 | 0.137 | 0.176 | 0.036 | 0.074 | 0.136 | 0.182 | 0.036 | 0.076 | 0.136 | |
| KOVAE-NCDE | 0.199 | 0.132 | 0.114 | 0.139 | 0.200 | 0.123 | 0.110 | 0.142 | 0.204 | 0.116 | 0.105 | 0.146 | |
| KOVAE | 0.170 | 0.094 | 0.080 | 0.139 | 0.172 | 0.076 | 0.080 | 0.140 | 0.175 | 0.061 | 0.080 | 0.144 | |
| GTGAN-NCDE | 0.216 | 0.106 | 0.116 | 0.183 | 0.209 | 0.100 | 0.122 | 0.186 | 0.199 | 0.108 | 0.123 | 0.193 | |
| GTGAN | 0.176 | 0.069 | 0.093 | 0.160 | 0.186 | 0.056 | 0.089 | 0.174 | 0.174 | 0.057 | 0.099 | 0.173 | |
Refined continuous TSG on solving for the analytical solutions of polynomial functions. We further assess continuous TSG using a cubic function . For each coefficient, 2500 samples are drawn from normal distributions with varying means and standard deviations. With linearly spaced over 24 values in , forms time series (visualized in Appendix figure 16). The model is trained on , aiming for coefficients derived from generated to match the original distributions.
Figure 4 shows that analytical solutions from our generated closely match the true distribution. Both irregular-to-regular and irregular-to-continuous TSG preserve distributional similarity, with refined continuous generation (Fig. 4(d)) closer to ground truth than non-continuous. This confirms the effectiveness of our method. In contrast, baselines using standard NCDE deviate significantly (Appendix Figures 17, 19, 20).
| 30% | 50% | 70% | ||||||||
| ECG200 | ECG5K | ECGFD | ECG200 | ECG5K | ECGFD | ECG200 | ECG5K | ECGFD | ||
| MDD | Ours | 0.211 | 0.062 | 0.424 | 0.237 | 0.079 | 0.507 | 0.246 | 0.095 | 0.624 |
| w/o MoE | 0.226 | 0.152 | 0.740 | 0.255 | 0.179 | 0.778 | 0.244 | 0.196 | 0.784 | |
| w/o Decoupled Design | 0.370 | 0.191 | 0.878 | 0.397 | 0.174 | 0.850 | 0.361 | 0.154 | 0.865 | |
| w/o MoE + w/o Decoupled Design | 0.412 | 0.173 | 0.965 | 0.363 | 0.175 | 0.968 | 0.361 | 0.157 | 0.984 | |
| KL | Ours | 0.052 | 0.009 | 0.055 | 0.107 | 0.014 | 0.037 | 0.098 | 0.02 | 0.037 |
| w/o MoE | 0.097 | 0.091 | 0.068 | 0.111 | 0.13 | 0.037 | 0.082 | 0.155 | 0.076 | |
| w/o Decoupled Design | 0.466 | 0.094 | 0.035 | 0.502 | 0.081 | 0.028 | 0.836 | 0.083 | 0.034 | |
| w/o MoE + w/o Decoupled Design | 0.877 | 0.109 | 1.203 | 0.583 | 0.096 | 1.211 | 0.667 | 0.074 | 1.323 | |
| 30% | 50% | 70% | ||||||||
| ECG200 | ECG5k | ECGFD | ECG200 | ECG5k | ECGFD | ECG200 | ECG5k | ECGFD | ||
| DS | Four-experts-eight-layers | 0.2 | 0.205 | 0.09 | 0.405 | 0.315 | 0.12 | 0.34 | 0.288 | 0.13 |
| One-expert-eight-layers | 0.328 | 0.364 | 0.13 | 0.44 | 0.389 | 0.23 | 0.392 | 0.412 | 0.21 | |
| MDD | Four-experts-eight-layers | 0.211 | 0.062 | 0.424 | 0.237 | 0.079 | 0.529 | 0.246 | 0.095 | 0.613 |
| One-expert-eight-layers | 0.288 | 0.12 | 0.694 | 0.321 | 0.152 | 0.831 | 0.314 | 0.165 | 0.837 | |
| KL | Four-experts-eight-layers | 0.052 | 0.009 | 0.055 | 0.107 | 0.014 | 0.05 | 0.098 | 0.02 | 0.049 |
| One-expert-eight-layers | 0.136 | 0.056 | 0.065 | 0.236 | 0.098 | 0.189 | 0.224 | 0.123 | 0.165 | |
4.4 Ablation study
Quantitative evaluation. Table 4 shows that removing MoE functions or the decoupled design significantly decreases model accuracy, and removing them together further deteriorates performance, indicating that these components are crucial. The MoE functions help NCDE learn temporal patterns, while the dynamics-focused decoupled optimization design (called Decoupled Design) uses a pre-trained channel-wise autoencoder to replace the network SIN and RN, allowing better focus on learning MoE dynamics.
Qualitative evaluation. We construct synthetic signals (e.g., sine, piecewise, sawtooth) and split them into training and test sets. We train MoE-NCDE and evaluate generation on the test set under component ablations (Fig. 5). Only with both MoE dynamics and the dynamics-focused decoupled optimization design does the model produce signals resembling the real data; removing either causes high errors, confirming their effectiveness.
Exploring why MoE is effective. Table 5 compares NCDE performance between (i) four experts (each with one MLP and two linear layers, eight layers total) and (ii) a single expert with four stacked MLPs (same eight layers). The MoE approach significantly outperforms the single expert, as each expert specializes in learning distinct temporal dynamics, making the overall dynamics easier to learn while enhancing both expressiveness and generalization. Figure 6 further illustrates how MoE decomposes complex temporal patterns across experts and then integrates them, substantially enhancing NCDE’s ability to model diverse temporal behaviors. Moreover, Appendix Table 13 reports the average expert weights across datasets. The weights are non-sparse, and different datasets favor different experts, indicating that the Mixture of Experts structure is effectively utilized rather than collapsing to a single expert.
4.5 Hyperparameter and time complexity analysis
Hyperparameter analysis. Figure 7 shows that MoE-NCDE performance is not highly sensitive to the number of experts: using more than one improves performance, with four experts yielding satisfactory results.
Time complexity analysis. MoE-NCDE has higher time complexity than standard NCDE (e.g., 1.2s vs. 0.81s for double-length refined generation on 3,674 stock samples of length 12) due to mixture-of-experts dynamics. However, by parameterizing the MoE weights via the diffusion model, we can directly generate sample-specific MoE weights for new data, avoiding NCDE retraining, additional feature extraction, and forward propagation, thereby reducing computational overhead.
5 Conclusion
This paper presents the first study of continuous TSG. Built upon NCDEs, Diff-MN introduces a decoupled MoE-NCDE that enhances standard NCDEs with mixture-of-experts dynamic functions and dynamics-focused optimization design, enabling more effective modeling of complex temporal dynamics and irregular TS for continuous generation. Moreover, we parameterize MoE weights via a diffusion model to generate sample-specific weights that reflect their temporal dynamics. Combined with a pre-trained NCDE, this enables adaptive continuous TSG. Diff-MN achieves state-of-the-art performance on continuous generation across multiple public and synthetic datasets.
Appendix A Supplementary related work
TSG for regular data.
TSG has gained increasing attention as a way to produce synthetic data for training, privacy preservation, and simulation. Early work explored adversarial learning approaches such as TimeGAN [30], which leverage adversarial objectives to jointly capture temporal dynamics and sample realism. Variational methods have also been developed, including TimeVAE [6] and Vector Quantized TimeVAE [12]. More recently, diffusion-based models [17] have emerged, synthesizing high-fidelity sequences by iteratively denoising random noise. These approaches illustrate the rapid methodological progress in generative modeling for TS. Recently, an increasing number of generation methods based on diffusion models have been proposed [4, 5, 9, 28, 14], such as task-oriented generation, cross-domain generation, causal generation, and text-controlled generation.
TS modeling with irregular data.
Irregularly sampled data becomes a critical challenge in TS modeling. An important line of work introduces continuous-time dynamics to naturally handle uneven sampling, such as Neural ODEs [2] and their extensions, which enrich the dynamics with additional inputs like the slope of imputed observations [11], also called Neural CDE. Other approaches directly encode timestamp information into sequence models, e.g., using specialized irregularity-aware architectures [31]. However, it cannot model the temporal patterns at arbitrary time points and cannot be directly applied to continuous generation tasks. Collectively, these methods underscore the importance of explicitly modeling temporal irregularity as a prerequisite for generating reliable and useful sequences.
Appendix B Discussion about why using neural controlled differential equation (NCDE)
B.1 Comparisons between NODE and NCDE
B.1.1 Theoretical perspective
Neural ordinary differential equation (NODE) [2] provides a continuous-time modeling framework for irregularly-sampled time series. The latent dynamics are governed by the following differential equation:
where is the latent state at time , and is a neural network parameterized by represent the dynamics of the differential equations.
However, NODE evolves the latent state solely from based on learned dynamics, without direct conditioning on the observed values. Consequently, the latent trajectory may drift away from the true data distribution.
We adopt NCDE because Diff-MN requires both imputing irregular observations and performing continuous generation. Values generated between existing observation points need to be consistent with the temporal dynamics of the surrounding context; otherwise, the generated values may deviate from the true distribution of the underlying dynamical system.
B.1.2 Experimental perspective
GT-GAN incorporates ODE-RNN, and ablation studies of GT-GAN consistently show that it outperforms the two ODE-based methods. Additionally, KoVAE (a TSG method using NCDE) is reported to perform even better than GT-GAN, and our Diff-MN outperforms both GT-GAN and KoVAE. All of this indicates that NCDE-based TSG under observed-data constraints outperforms ODE-based methods.
B.2 Proving interpolation control in NCDE does not cause over-smoothing
B.2.1 Theoretical perspective
As discussed in the main text, NCDE can be defined as
where is the control path (obtained via interpolation and smoothing), and is a nonlinear function (neural network).
The neural network can model highly nonlinear dynamics, amplifying or bending small variations in the smooth control path and thereby inducing complex, rapidly changing behavior in the latent trajectory.
Hence, even if is smooth, the behavior of the integral result depends on the properties of . If is highly nonlinear, the derivative can vary rapidly with , leading to non-smooth trajectories.
Specifically, when is smooth and differentiable, the NCDE admits the following differential form:
Although the interpolated control path is smooth by construction, the evolution of is governed by the nonlinear function . As a result, even when is smooth, the product may vary rapidly due to the nonlinearity of , leading to non-smooth latent trajectories.
In our model, the is further parameterized as a Mixture-of-Experts:
where and are routing weights produced by the router , and denotes the -th expert.
Substituting the MoE formulation into the NCDE dynamics yields
Since both the expert functions and the routing weights depend nonlinearly on the latent state, different experts may dominate the dynamics at different time instances. This allows the latent trajectory to exhibit rapid local variations and highly diverse temporal behaviors.
Therefore, even with a smooth interpolated control path , the NCDE equipped with MoE dynamics can generate expressive and potentially non-smooth latent trajectories.
B.2.2 Experimental perspective
To validate this, we added visualizations of sequences generated solely by NCDE, as shown Figure 13, 14 and 15. These results show: (i) For intrinsically smooth data (e.g., sine in Figure 16), NCDE produces smooth trajectories and (ii) For intrinsically smooth data (e.g., sine in Figure 16), NCDE produces smooth trajectories.
Overall, while the control path is smooth, the nonlinear MoE–NCDE dynamics allow the generated sequences to exhibit fluctuating temporal behaviors, depending entirely on the underlying temporal dynamics.
Appendix C Supplementary algorithm pseudocodes
This presents the algorithm pseudocodes omitted from the Method section of the main text due to space constraints. As all algorithms have already been properly referenced and discussed, we simply include them here for convenient reference.
C.1 Decoupled optimization for MoE-dynamics-focused training (Algorithm 2)
C.2 Training channel-wise autoencoder (Algorithm 1)
C.3 Joint diffusion training for to parameterize MoE weights (Algorithm 3)
C.4 Continuous time series generation (Algorithm 4)
Appendix D Supplementary experimental settings
D.1 Datasets
We evaluate on four popular TSG datasets with diverse characteristics, following established benchmarks for generative time series [30, 10]. Sines is a simulated multivariate dataset where each channel follows a sinusoidal function with random frequency and phase, capturing continuous and periodic patterns. Stocks contains daily Google stock data (2004–2019) across six financial indicators, exhibiting aperiodic, random-walk behaviors. Energy is the UCI appliance energy prediction dataset [1], comprising 28 correlated channels with noisy periodicity and continuous-valued measurements. Finally, MuJoCo (Multi-Joint dynamics with Contact) [25] provides simulated physical dynamics with 14 channels.
The used medical ECG datasets are ECG200, ECG5000 (ECG5K), ECGFiveDays (ECGFD), and TwoLeadECG (TLECG), which can be downloaded from the link111www.timeseriesclassification.com. Their sequence lengths are 96, 140, 136, and 82. The number of classes is 2, 5, 2, and 2, respectively. The data sizes are 200, 5000, 884, and 1162, respectively.
D.2 More implementation details
When training MoE-NCDE, we use the Adam optimizer with a learning rate of , a batch size of 256, and a fixed hidden dimension of 64. For the TSG backbone of the standard diffusion model, the denoising network adopts a U-Net with four down- and up-sampling stages. Each stage contains two residual blocks and one cross-attention block, where residual blocks use two 1D convolution layers, and cross-attention employs 1D convolutions for input/output projections with eight attention heads. A middle block, placed between the down- and up-sampling paths, consists of two residual blocks and one attention block. In addition, we include input and output layers, each implemented with a single 1D convolution. SiLU is used as the activation function throughout. We train the model for 600 epochs with a batch size of 256, a learning rate of .
Except for the sequence length, all other dataset settings follow previous work [19]. More implementation details can be found in our code link222https://github.com/microsoft/TimeCraft/tree/main/Diff-MN. All model parameters used in the experiments are specified directly in the code.
Appendix E Supplementary experimental results
In Table 1 of the main text, we report the average performance of irregular-to-regular generation tasks across sequence lengths of 12, 24, and 36. Here, we additionally present results for each individual length. Overall, our method consistently outperforms others, in line with the conclusions from Table 1, and we therefore omit further discussion. We simply include them here for convenient reference.
E.1 Irregular to regular TSG on popular datasets with sequence length 12 (Table 6)
E.2 Irregular to regular TSG on popular datasets with sequence length 24 (Table 7)
E.3 Irregular to regular TSG on popular datasets with sequence length 36 (Table 8)
E.4 TSNE and PDF visualizations on irregular to regular TSG (Figure 9)
E.5 Evaluation of privacy protection capability (Table 9)
E.6 Visualizations about MoE-NCDE not producing overly smooth curves (Figure 13, 14 and 15)
E.7 Irregular to continuous TSG on popular datasets with sequence length 12 (Table 10)
E.8 Irregular to continuous TSG on popular datasets with sequence length 24 (Table 11)
E.9 Irregular to continuous TSG on popular datasets with sequence length 36 (Table 12)
E.10 Visualization of continuous TSG on the Stock dataset (Figure 10 for part 1, Figure 11 for part 2, and Figure 12 for part 3)
E.11 Visualization of cubic polynomial curves (Figure 16)
E.12 Cubic polynomial solutions of Diff-MN with 50% and 70% missing values (Figure 17)
E.13 Cubic polynomial solutions of Diff-MN VS. KOVAE with 30%, 50% and 70% missing values (Figure 19)
E.14 Cubic polynomial solutions of Diff-MN VS. GT-GAN with 30%, 50% and 70% missing values (Figure 20)
E.15 Visualization of temporal dynamics learned by different expert (Figure 18)
E.16 Average expert weights learned by MoE-NCDE (Figure 18)
E.17 The impact of the number of experts on the DS metric (Figure 8)
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| DS | Ours | 0.077 | 0.084 | 0.386 | 0.268 | 0.13 | 0.042 | 0.472 | 0.417 | 0.222 | 0.054 | 0.494 | 0.45 |
| KoVAE | 0.218 | 0.103 | 0.500 | 0.358 | 0.202 | 0.106 | 0.500 | 0.390 | 0.211 | 0.228 | 0.500 | 0.451 | |
| GT-GAN | 0.300 | 0.385 | 0.499 | 0.498 | 0.423 | 0.357 | 0.499 | 0.496 | 0.434 | 0.370 | 0.498 | 0.497 | |
| TimeGAN-NCDE | 0.468 | 0.369 | 0.498 | 0.487 | 0.422 | 0.472 | 0.499 | 0.499 | 0.429 | 0.479 | 0.500 | 0.498 | |
| TimeVAE-NCDE | 0.354 | 0.493 | 0.496 | 0.433 | 0.466 | 0.497 | 0.491 | 0.477 | 0.500 | 0.497 | 0.498 | 0.480 | |
| Diffusion-NCDE | 0.264 | 0.428 | 0.426 | 0.323 | 0.353 | 0.472 | 0.482 | 0.437 | 0.480 | 0.491 | 0.497 | 0.475 | |
| ProFITi | 0.29 | 0.44 | 0.492 | 0.49 | 0.339 | 0.478 | 0.498 | 0.496 | 0.41 | 0.486 | 0.5 | 0.497 | |
| HeTVAE | 0.5 | 0.368 | 0.493 | 0.499 | 0.5 | 0.478 | 0.499 | 0.5 | 0.5 | 0.493 | 0.5 | 0.5 | |
| MDD | Ours | 1.015 | 0.16 | 0.303 | 0.303 | 1.337 | 0.2 | 0.251 | 0.319 | 1.783 | 0.219 | 0.319 | 0.34 |
| KoVAE | 10.137 | 0.289 | 0.343 | 0.326 | 10.12 | 0.288 | 0.377 | 0.372 | 10.13 | 0.312 | 0.399 | 0.635 | |
| GT-GAN | 3.574 | 0.746 | 0.598 | 0.540 | 3.898 | 0.600 | 0.414 | 0.565 | 3.38 | 0.733 | 0.562 | 0.585 | |
| TimeGAN-NCDE | 4.599 | 1.109 | 1.020 | 1.261 | 4.799 | 1.625 | 1.395 | 1.485 | 5.35 | 1.717 | 1.499 | 1.761 | |
| TimeVAE-NCDE | 8.661 | 0.712 | 0.451 | 0.463 | 6.045 | 0.864 | 0.587 | 0.586 | 5.184 | 1.054 | 0.652 | 0.707 | |
| Diffusion-NCDE | 2.150 | 0.328 | 0.306 | 0.276 | 2.128 | 0.538 | 0.365 | 0.428 | 2.902 | 0.854 | 0.506 | 0.508 | |
| ProFITi | 1.277 | 0.32 | 0.254 | 0.298 | 1.342 | 0.416 | 0.342 | 0.361 | 1.584 | 0.443 | 0.442 | 0.446 | |
| HeTVAE | 3.594 | 0.597 | 0.635 | 0.507 | 4.25 | 0.748 | 1.11 | 1.092 | 5.134 | 0.986 | 1.273 | 1.105 | |
| KL | Ours | 0.007 | 0.05 | 0.019 | 0.017 | 0.014 | 0.075 | 0.025 | 0.018 | 0.033 | 0.066 | 0.013 | 0.024 |
| KoVAE | 9.941 | 0.139 | 0.076 | 0.168 | 9.511 | 0.149 | 0.084 | 0.282 | 9.788 | 0.151 | 0.111 | 0.579 | |
| GT-GAN | 0.144 | 0.283 | 0.063 | 0.020 | 0.187 | 0.212 | 0.050 | 0.029 | 0.164 | 0.314 | 0.066 | 0.031 | |
| TimeGAN-NCDE | 5.230 | 1.742 | 0.210 | 0.575 | 1.125 | 7.401 | 0.716 | 0.997 | 15.176 | 12.148 | 0.854 | 2.397 | |
| TimeVAE-NCDE | 3.551 | 0.298 | 0.089 | 0.158 | 2.776 | 0.345 | 0.128 | 0.245 | 2.113 | 0.534 | 0.148 | 0.338 | |
| Diffusion-NCDE | 0.004 | 0.012 | 0.004 | 0.007 | 0.028 | 0.040 | 0.019 | 0.036 | 0.074 | 0.095 | 0.046 | 0.042 | |
| ProFITi | 0.078 | 0.084 | 0.022 | 0.016 | 0.082 | 0.124 | 0.032 | 0.016 | 0.11 | 0.137 | 0.044 | 0.032 | |
| HeTVAE | 0.036 | 0.302 | 0.376 | 0.102 | 2.655 | 0.46 | 1.124 | 0.934 | 0.288 | 0.78 | 1.275 | 1.079 | |
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| DS | Ours | 0.069 | 0.171 | 0.414 | 0.249 | 0.072 | 0.151 | 0.491 | 0.306 | 0.127 | 0.179 | 0.499 | 0.347 |
| KoVAE | 0.055 | 0.285 | 0.500 | 0.358 | 0.090 | 0.208 | 0.500 | 0.363 | 0.212 | 0.236 | 0.500 | 0.403 | |
| GT-GAN | 0.276 | 0.305 | 0.500 | 0.480 | 0.338 | 0.325 | 0.500 | 0.491 | 0.286 | 0.273 | 0.500 | 0.489 | |
| TimeGAN-NCDE | 0.457 | 0.351 | 0.500 | 0.486 | 0.433 | 0.454 | 0.500 | 0.497 | 0.438 | 0.472 | 0.499 | 0.499 | |
| TimeVAE-NCDE | 0.191 | 0.481 | 0.499 | 0.324 | 0.175 | 0.493 | 0.499 | 0.462 | 0.431 | 0.490 | 0.499 | 0.464 | |
| Diffusion-NCDE | 0.148 | 0.460 | 0.455 | 0.373 | 0.286 | 0.490 | 0.489 | 0.409 | 0.392 | 0.489 | 0.496 | 0.470 | |
| ProFITi | 0.298 | 0.459 | 0.498 | 0.496 | 0.365 | 0.487 | 0.499 | 0.498 | 0.409 | 0.492 | 0.5 | 0.498 | |
| HeTVAE | 0.48 | 0.306 | 0.498 | 0.499 | 0.5 | 0.401 | 0.5 | 0.499 | 0.5 | 0.476 | 0.5 | 0.5 | |
| MDD | Ours | 0.752 | 0.269 | 0.218 | 0.336 | 0.775 | 0.251 | 0.216 | 0.341 | 0.971 | 0.299 | 0.265 | 0.285 |
| KoVAE | 2.756 | 0.905 | 0.347 | 0.341 | 3.314 | 0.833 | 0.361 | 0.380 | 5.221 | 0.763 | 0.368 | 0.633 | |
| GT-GAN | 2.257 | 0.551 | 0.429 | 0.576 | 2.467 | 0.563 | 0.396 | 0.532 | 2.277 | 0.532 | 0.413 | 0.555 | |
| TimeGAN-NCDE | 3.162 | 1.134 | 1.046 | 1.173 | 3.555 | 1.597 | 1.226 | 1.553 | 3.609 | 1.592 | 1.605 | 1.651 | |
| TimeVAE-NCDE | 3.408 | 0.580 | 0.418 | 0.307 | 3.042 | 0.709 | 0.521 | 0.411 | 2.862 | 0.804 | 0.597 | 0.527 | |
| Diffusion-NCDE | 1.303 | 0.401 | 0.250 | 0.336 | 1.402 | 0.403 | 0.356 | 0.338 | 1.667 | 0.692 | 0.557 | 0.486 | |
| ProFITi | 0.961 | 0.311 | 0.261 | 0.284 | 1.446 | 0.472 | 0.358 | 0.333 | 1.645 | 0.436 | 0.444 | 0.407 | |
| HeTVAE | 2.775 | 0.582 | 0.607 | 0.722 | 3.431 | 0.787 | 0.993 | 1.09 | 3.677 | 1.071 | 1.229 | 1.445 | |
| KL | Ours | 0.011 | 0.084 | 0.009 | 0.007 | 0.022 | 0.092 | 0.014 | 0.006 | 0.035 | 0.105 | 0.023 | 0.013 |
| KoVAE | 1.696 | 1.347 | 0.058 | 0.150 | 1.861 | 1.205 | 0.061 | 0.198 | 2.840 | 1.380 | 0.071 | 0.508 | |
| GT-GAN | 0.065 | 0.163 | 0.042 | 0.032 | 0.139 | 0.161 | 0.053 | 0.019 | 0.190 | 0.184 | 0.051 | 0.034 | |
| TimeGAN-NCDE | 4.573 | 1.380 | 0.227 | 0.666 | 4.339 | 5.887 | 0.489 | 1.248 | 4.854 | 10.508 | 1.119 | 1.409 | |
| TimeVAE-NCDE | 2.18 | 0.209 | 0.067 | 0.046 | 1.291 | 0.245 | 0.102 | 0.070 | 1.038 | 0.284 | 0.132 | 0.112 | |
| Diffusion-NCDE | 0.060 | 0.438 | 0.073 | 0.033 | 0.053 | 0.440 | 0.099 | 0.041 | 0.077 | 0.498 | 0.114 | 0.047 | |
| ProFITi | 0.128 | 0.091 | 0.028 | 0.01 | 0.137 | 0.161 | 0.036 | 0.012 | 0.189 | 0.123 | 0.043 | 0.026 | |
| HeTVAE | 0.323 | 0.792 | 0.175 | 0.318 | 8.712 | 1.203 | 0.495 | 0.915 | 0.337 | 1.398 | 1.429 | 2.343 | |
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| DS | Ours | 0.169 | 0.170 | 0.466 | 0.363 | 0.183 | 0.218 | 0.499 | 0.403 | 0.196 | 0.085 | 0.499 | 0.381 |
| KoVAE | 0.152 | 0.286 | 0.429 | 0.399 | 0.222 | 0.247 | 0.500 | 0.439 | 0.261 | 0.176 | 0.500 | 0.422 | |
| GT-GAN | 0.331 | 0.272 | 0.499 | 0.489 | 0.487 | 0.297 | 0.500 | 0.494 | 0.303 | 0.234 | 0.500 | 0.489 | |
| TimeGAN-NCDE | 0.440 | 0.376 | 0.499 | 0.496 | 0.425 | 0.399 | 0.498 | 0.500 | 0.453 | 0.48 | 0.500 | 0.499 | |
| TimeVAE-NCDE | 0.157 | 0.437 | 0.500 | 0.368 | 0.282 | 0.464 | 0.499 | 0.451 | 0.347 | 0.490 | 0.498 | 0.485 | |
| Diffusion-NCDE | 0.189 | 0.435 | 0.498 | 0.368 | 0.352 | 0.462 | 0.496 | 0.436 | 0.391 | 0.493 | 0.497 | 0.474 | |
| ProFITi | 0.342 | 0.481 | 0.499 | 0.498 | 0.354 | 0.493 | 0.5 | 0.498 | 0.357 | 0.494 | 0.5 | 0.498 | |
| HeTVAE | 0.494 | 0.309 | 0.499 | 0.498 | 0.5 | 0.449 | 0.5 | 0.499 | 0.5 | 0.493 | 0.5 | 0.5 | |
| MDD | Ours | 1.091 | 0.322 | 0.288 | 0.402 | 1.167 | 0.392 | 0.288 | 0.293 | 1.17 | 0.38 | 0.253 | 0.267 |
| KoVAE | 2.509 | 0.933 | 0.350 | 0.456 | 3.319 | 0.865 | 0.366 | 0.506 | 3.828 | 0.798 | 0.363 | 0.588 | |
| GT-GAN | 2.101 | 0.593 | 0.673 | 0.558 | 2.568 | 0.604 | 0.610 | 0.692 | 2.021 | 0.543 | 0.620 | 0.74 | |
| TimeGAN-NCDE | 2.670 | 1.091 | 1.019 | 1.111 | 2.863 | 1.592 | 1.408 | 1.568 | 2.984 | 1.578 | 1.574 | 1.7 | |
| TimeVAE-NCDE | 2.232 | 0.499 | 0.463 | 0.243 | 2.195 | 0.604 | 0.491 | 0.331 | 2.144 | 0.670 | 0.564 | 0.462 | |
| Diffusion-NCDE | 0.934 | 0.295 | 0.319 | 0.308 | 0.963 | 0.487 | 0.331 | 0.345 | 1.421 | 0.682 | 0.475 | 0.468 | |
| ProFITi | 0.935 | 0.252 | 0.259 | 0.289 | 1.024 | 0.394 | 0.366 | 0.337 | 1.161 | 0.45 | 0.444 | 0.452 | |
| HeTVAE | 3.104 | 0.619 | 0.634 | 0.734 | 2.665 | 0.784 | 1.182 | 1.098 | 3.091 | 1.038 | 1.466 | 1.644 | |
| KL | Ours | 0.020 | 0.089 | 0.031 | 0.039 | 0.033 | 0.115 | 0.026 | 0.003 | 0.031 | 0.101 | 0.014 | 0.004 |
| KoVAE | 0.878 | 1.555 | 0.061 | 0.072 | 2.785 | 1.224 | 0.057 | 0.115 | 3.935 | 1.072 | 0.057 | 0.188 | |
| GT-GAN | 0.087 | 0.163 | 0.064 | 0.012 | 0.134 | 0.171 | 0.051 | 0.035 | 0.134 | 0.171 | 0.054 | 0.078 | |
| TimeGAN-NCDE | 5.086 | 2.104 | 0.341 | 0.440 | 3.398 | 7.039 | 0.551 | 1.214 | 17.403 | 12.235 | 1.012 | 1.625 | |
| TimeVAE-NCDE | 0.265 | 0.166 | 0.107 | 0.021 | 0.395 | 0.186 | 0.074 | 0.038 | 0.437 | 0.216 | 0.128 | 0.067 | |
| Diffusion-NCDE | 0.087 | 0.235 | 0.151 | 0.042 | 0.083 | 0.334 | 0.044 | 0.071 | 0.123 | 0.398 | 0.163 | 0.039 | |
| ProFITi | 0.08 | 0.054 | 0.022 | 0.012 | 0.172 | 0.115 | 0.036 | 0.011 | 0.192 | 0.121 | 0.047 | 0.029 | |
| HeTVAE | 3.174 | 2.16 | 0.194 | 0.353 | 1.34 | 2.537 | 1.131 | 1.14 | 6.838 | 2.659 | 2.03 | 3.184 | |
| 30% | 50% | 70% | ||||||||||||
| ECG200 | ECG5k | ECGFD | TLECG | ECG200 | ECG5k | ECGFD | TLECG | ECG200 | ECG5k | ECGFD | TLECG | Avg. | ||
| MIR | Ours | 0.6826 | 0.6944 | 0.6734 | 0.6671 | 0.6671 | 0.6676 | 0.6866 | 0.7302 | 0.697 | 0.6866 | 0.6765 | 0.6866 | 0.6846 |
| KoVAE | 0.6873 | 0.6849 | 0.7194 | 0.6882 | 0.6761 | 0.6873 | 0.8846 | 0.8679 | 0.7931 | 0.7077 | 0.697 | 0.697 | 0.7325 | |
| GT-GAN | 0.6993 | 0.8772 | 0.9709 | 0.9794 | 0.7698 | 0.7628 | 0.9583 | 0.9787 | 0.6765 | 0.7541 | 0.807 | 0.7667 | 0.8334 | |
| TimeGAN-NCDE | 0.7463 | 0.7463 | 0.7463 | 0.6817 | 0.7163 | 0.8482 | 0.807 | 0.807 | 0.807 | 0.7667 | 0.7419 | 0.807 | 0.7685 | |
| TimeVAE-NCDE | 0.7194 | 0.7067 | 0.6897 | 0.6676 | 0.6676 | 0.6698 | 0.7541 | 0.7541 | 0.7797 | 0.7419 | 0.7419 | 0.7419 | 0.7195 | |
| Diffusion-NCDE | 0.7067 | 0.7067 | 0.6873 | 0.6693 | 0.6698 | 0.6698 | 0.7541 | 0.7797 | 0.8364 | 0.7419 | 0.7188 | 0.7302 | 0.7226 | |
| ProFITi | 0.7067 | 0.7042 | 0.7194 | 0.6988 | 0.6798 | 0.6969 | 0.7797 | 0.902 | 0.807 | 0.7797 | 0.7188 | 0.8846 | 0.7565 | |
| HetVAE | 0.7692 | 0.6993 | 0.7117 | 0.7981 | 0.8163 | 0.8299 | 0.7931 | 0.8214 | 0.807 | 0.7188 | 0.7302 | 0.8519 | 0.7789 | |
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| MSE | Ours- | 0.016 | 0.002 | 0.014 | 0.027 | 0.017 | 0.002 | 0.013 | 0.029 | 0.017 | 0.002 | 0.014 | 0.027 |
| Ours | 0.030 | 0.004 | 0.014 | 0.025 | 0.031 | 0.005 | 0.014 | 0.027 | 0.034 | 0.004 | 0.014 | 0.026 | |
| KOVAE-NCDE | 0.070 | 0.013 | 0.03 | 0.028 | 0.070 | 0.012 | 0.028 | 0.029 | 0.07 | 0.011 | 0.027 | 0.031 | |
| KOVAE | 0.036 | 0.006 | 0.014 | 0.027 | 0.036 | 0.006 | 0.014 | 0.027 | 0.036 | 0.006 | 0.014 | 0.028 | |
| GTGAN-NCDE | 0.064 | 0.039 | 0.029 | 0.041 | 0.052 | 0.028 | 0.024 | 0.040 | 0.060 | 0.028 | 0.030 | 0.040 | |
| GTGAN | 0.034 | 0.029 | 0.018 | 0.032 | 0.043 | 0.011 | 0.016 | 0.031 | 0.033 | 0.026 | 0.018 | 0.030 | |
| MAE | Ours- | 0.107 | 0.025 | 0.077 | 0.124 | 0.109 | 0.025 | 0.074 | 0.127 | 0.111 | 0.024 | 0.076 | 0.126 |
| Ours | 0.140 | 0.034 | 0.077 | 0.124 | 0.143 | 0.035 | 0.075 | 0.128 | 0.149 | 0.033 | 0.077 | 0.127 | |
| KOVAE-NCDE | 0.190 | 0.085 | 0.137 | 0.132 | 0.190 | 0.083 | 0.134 | 0.134 | 0.190 | 0.079 | 0.128 | 0.141 | |
| KOVAE | 0.157 | 0.039 | 0.080 | 0.131 | 0.157 | 0.040 | 0.080 | 0.130 | 0.156 | 0.041 | 0.080 | 0.135 | |
| GTGAN-NCDE | 0.193 | 0.147 | 0.111 | 0.162 | 0.181 | 0.122 | 0.115 | 0.162 | 0.183 | 0.122 | 0.121 | 0.16 | |
| GTGAN | 0.153 | 0.110 | 0.090 | 0.141 | 0.174 | 0.077 | 0.087 | 0.142 | 0.15 | 0.103 | 0.093 | 0.139 | |
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| MSE | Ours- | 0.02 | 0.003 | 0.013 | 0.027 | 0.020 | 0.002 | 0.013 | 0.027 | 0.022 | 0.002 | 0.013 | 0.027 |
| Ours | 0.055 | 0.007 | 0.013 | 0.030 | 0.055 | 0.006 | 0.013 | 0.029 | 0.059 | 0.006 | 0.014 | 0.029 | |
| KOVAE-NCDE | 0.080 | 0.053 | 0.021 | 0.030 | 0.080 | 0.049 | 0.018 | 0.032 | 0.084 | 0.039 | 0.018 | 0.033 | |
| KOVAE | 0.044 | 0.033 | 0.014 | 0.029 | 0.044 | 0.022 | 0.014 | 0.030 | 0.047 | 0.018 | 0.014 | 0.031 | |
| GTGAN-NCDE | 0.077 | 0.010 | 0.02 | 0.056 | 0.079 | 0.01 | 0.023 | 0.053 | 0.073 | 0.013 | 0.022 | 0.058 | |
| GTGAN | 0.047 | 0.007 | 0.017 | 0.038 | 0.047 | 0.005 | 0.016 | 0.048 | 0.048 | 0.003 | 0.016 | 0.047 | |
| MAE | Ours- | 0.106 | 0.025 | 0.074 | 0.131 | 0.107 | 0.025 | 0.072 | 0.129 | 0.109 | 0.025 | 0.074 | 0.131 |
| Ours | 0.179 | 0.040 | 0.074 | 0.138 | 0.18 | 0.037 | 0.073 | 0.137 | 0.187 | 0.038 | 0.075 | 0.135 | |
| KOVAE-NCDE | 0.191 | 0.152 | 0.109 | 0.140 | 0.192 | 0.144 | 0.096 | 0.143 | 0.199 | 0.135 | 0.096 | 0.146 | |
| KOVAE | 0.166 | 0.119 | 0.08 | 0.137 | 0.167 | 0.079 | 0.08 | 0.139 | 0.176 | 0.071 | 0.080 | 0.143 | |
| GTGAN-NCDE | 0.19 | 0.079 | 0.104 | 0.188 | 0.199 | 0.078 | 0.115 | 0.188 | 0.197 | 0.086 | 0.108 | 0.194 | |
| GTGAN | 0.179 | 0.052 | 0.087 | 0.157 | 0.173 | 0.045 | 0.084 | 0.177 | 0.173 | 0.034 | 0.086 | 0.173 | |
| 30% | 50% | 70% | |||||||||||
| Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | Sines | Stocks | Energy | MuJoCo | ||
| MSE | Ours- | 0.041 | 0.002 | 0.013 | 0.031 | 0.042 | 0.002 | 0.013 | 0.031 | 0.044 | 0.003 | 0.013 | 0.032 |
| Ours-NCDE | 0.041 | 0.002 | 0.013 | 0.031 | 0.042 | 0.002 | 0.013 | 0.031 | 0.044 | 0.003 | 0.013 | 0.033 | |
| Ours | 0.063 | 0.005 | 0.013 | 0.038 | 0.064 | 0.005 | 0.014 | 0.035 | 0.068 | 0.005 | 0.014 | 0.036 | |
| KOVAE-NCDE | 0.077 | 0.057 | 0.018 | 0.033 | 0.078 | 0.046 | 0.018 | 0.035 | 0.082 | 0.039 | 0.017 | 0.036 | |
| KOVAE | 0.051 | 0.035 | 0.014 | 0.035 | 0.052 | 0.029 | 0.014 | 0.036 | 0.051 | 0.018 | 0.014 | 0.038 | |
| GTGAN-NCDE | 0.126 | 0.013 | 0.032 | 0.061 | 0.098 | 0.015 | 0.031 | 0.065 | 0.076 | 0.021 | 0.034 | 0.076 | |
| GTGAN | 0.052 | 0.006 | 0.022 | 0.054 | 0.064 | 0.007 | 0.020 | 0.062 | 0.058 | 0.004 | 0.023 | 0.065 | |
| MAE | Ours- | 0.162 | 0.026 | 0.074 | 0.133 | 0.161 | 0.027 | 0.074 | 0.135 | 0.163 | 0.026 | 0.075 | 0.137 |
| Ours | 0.203 | 0.035 | 0.074 | 0.149 | 0.204 | 0.036 | 0.074 | 0.144 | 0.209 | 0.036 | 0.076 | 0.145 | |
| KOVAE-NCDE | 0.214 | 0.159 | 0.096 | 0.146 | 0.218 | 0.141 | 0.099 | 0.149 | 0.225 | 0.133 | 0.091 | 0.151 | |
| KOVAE | 0.186 | 0.123 | 0.080 | 0.150 | 0.191 | 0.111 | 0.081 | 0.151 | 0.193 | 0.072 | 0.081 | 0.155 | |
| GTGAN-NCDE | 0.266 | 0.093 | 0.132 | 0.198 | 0.247 | 0.099 | 0.135 | 0.208 | 0.218 | 0.115 | 0.141 | 0.226 | |
| GTGAN | 0.197 | 0.046 | 0.103 | 0.182 | 0.209 | 0.048 | 0.098 | 0.202 | 0.201 | 0.034 | 0.117 | 0.208 | |
| Datasets | Expert 1 | Expert 2 | Expert 3 | Expert 4 | |
| Sines | 0.31384414 | 0.2231202 | 0.26881662 | 0.19421977 | |
| Stocks | 0.19200034 | 0.21767974 | 0.32304975 | 0.2672697 | |
| Energy | 0.2413441 | 0.31066164 | 0.2065838 | 0.24141027 | |
| MuJoCo | 0.18345672 | 0.241069 | 0.2046078 | 0.370866 | |
| ECG200 | 0.21229412 | 0.44273788 | 0.20304906 | 0.14191894 | |
| ECG5K | 0.21999635 | 0.182932 | 0.3610456 | 0.2360259 | |
| ECGFD | 0.17836495 | 0.31938562 | 0.2739047 | 0.2283447 | |
| TLECG | 0.19264506 | 0.2577268 | 0.3210036 | 0.2286246 |